seed cost .20/oz
nuts cost .30/oz
a 42oz mixture costs $11.00
how many ounces of seeds and nuts were used in this mixture?
I do not know how to put this into an algebra problem. Thank you.

Answer by stanbon(75887)

seed cost .20/oz
nuts cost .30/oz
a 42oz mixture costs $11.00
how many ounces of seeds and nuts were used in this mixture?
---------------
Equations:
Weight Eq.::: s + n = 42 oz
Value Eq. :: 20s + 30n = 1100 cents
------------------------------------------
Multiply the 1st equation by 20 to get:
20s + 20n = 20*42
-----
Subtract that from the Value Eq. and solve for "n":
10n = 1100-840
10n = 260
n = 26 oz (# of ounces of nuts needed in the mix)
----
since s+n = 42, s = 16 oz (# of ounces of seeds needed in the mix)
=============================================
Cheers,
Stan H.